# mode = titleUSAD

### v1

@@ 2.1 Deep Feature Extraction
High-quality feature extraction is fundamental to the accurate classification and recognition of complex human activities. In recent years, research focusing on neural network architecture design and the integration of multi-source HAR features has seen significant advancements. However, conventional ResNet architectures still suffer from weak inter-channel correlations, parameter redundancy, and insufficient feature reutilization. To address these limitations, Zhang et al. [1] proposed the ResNeSt network, which introduces a Split-Attention module and a restructured feature aggregation mechanism to enhance the modeling of channel dependencies, thereby further improving feature representation quality. Mekruksavanich et al. [2] adopted the ResNeXt architecture, integrating both inertial and stretch sensor data to achieve efficient human activity classification. Other studies have enriched feature extraction methodologies from various perspectives. Y. Wang et al. [3] introduced the innovative DMEFAM framework, which integrates temporal, channel, and spatial feature representations by leveraging components such as Bi-GRU, self-attention mechanisms, CBAM, and ResNet-18. This approach not only improved recognition accuracy but also demonstrated robust generalization capability on the DAAD dataset. Zhang J. et al. [4] employed a contrastive learning mechanism to extract activity style features from multi-source data with high precision. These features preserve both data characteristics and class semantic information, playing a critical role in multi-source feature fusion and effectively enhancing cross-subject generalization, thus offering a novel solution to domain adaptation challenges. Liu X. et al. [5] proposed a new multi-source feature fusion method for HAR, wherein time series data are encoded into RGB images through MTF, RP, and GAF techniques to capture multidimensional features. They further designed CIR-DFENet, which processes these images using multi-scale convolution and global attention mechanisms, and combines this with a 1D CNN+LSTM structure and self-attention module, thereby enabling effective fusion of multi-source features.
@@ 2.2 Data Augmentation
A variety of data augmentation methods have been proposed in previous studies to address issues of data scarcity and model overfitting. In earlier work, J. Wang [6] introduced SensoryGANs, attempting to generate synthetic sensor data using generative adversarial networks (GANs) to improve activity recognition accuracy, and also proposed visualization methods for evaluating the quality of synthetic data. However, SensoryGANs face challenges such as unstable training processes and the need to train separate models for each activity, thereby reducing efficiency. Li, X [7] proposed ActivityGAN, which aimed to generate multi-activity data within a unified framework. ActivityGAN can effectively generate sensor data that simulate human activities, expanding the dataset and improving model training and classification accuracy. Nevertheless, it did not achieve a substantial improvement in recognition performance and exhibited significant shortcomings in cross-activity generalization ability.
To overcome these limitations of traditional approaches, some researchers have turned to denoising diffusion probabilistic models (DDPM) [8], a class of generative models grounded in non-equilibrium thermodynamics, capable of producing high-quality images through unique training and sampling mechanisms. S. Shao [9] designed a diffusion modeling framework specifically adapted to sensor data characteristics based on DDPM. This study reconstructs the UNet architecture, adjusts the input layer to accommodate low-channel and varying sequence-length data, streamlines the sampling layers to avoid over-feature extraction, and embeds class labels for conditional generation. On the HASC and PAMAP2 datasets, the quality of generated data surpassed traditional transformations and GAN-based methods; blending real and synthetic data (at a ratio of 1:4) further improved model performance. Existing diffusion model-based methods are primarily designed for unimodal data (e.g., inertial sensors). In HAR tasks, research on the collaborative augmentation of multimodal data (such as IMU sensor data) remains inadequate. Although Zhou, Y [10] employed gradient optimization to determine the probability distribution of augmentation strategies and achieved automatic and efficient augmentation policy search—operating in two stages, accommodating multimodal properties, being model-agnostic and highly scalable, and significantly enhancing HAR data augmentation—its performance gains in adjacent domains (e.g., EEG classification) were limited, reflecting that modality specificity imposes constraints on the universality of augmentation strategies.
@@ 2.3 Attention Mechanisms
In recent years, deep learning models based on attention mechanisms have garnered significant attention. Abedini, A. [11] proposed the Att-CNN-BLSTM model, which employs CNNs and BLSTM units to extract spatial features and analyze temporal patterns, respectively, and introduces an attention layer to enhance the discriminability of features. This model achieved a classification accuracy of 90.54% and an F1 score of 86.88%, outperforming traditional models. However, its attention mechanism suffers from high computational complexity and limited interpretability. Ye, X. et al. [12] proposed the Deep Generative Domain Adaptation with Temporal Attention (DGDA) method, which leverages generative models in conjunction with temporal relational attention mechanisms to effectively exploit temporal dependencies within time series data, thereby improving distribution alignment and enhancing classification performance for HAR tasks. Experimental results on multiple public sensor HAR datasets indicate that DGDA exhibits excellent cross-user performance. Nevertheless, its capability to fuse heterogeneous sensor data in complex, real-world scenarios remains limited, and the high computational overhead poses challenges for deployment on resource-constrained devices. Zhao, J. et al. [13] introduced a DDoS attack detection method combining self-attention mechanisms with CNN-BiLSTM architecture. By concurrently extracting spatial and temporal sequence features using 1D CNNs and BiLSTM, and employing attention mechanisms to emphasize critical features, their method achieved an accuracy of up to 95.67% on the CIC-ISD 2017 and CIC-DDoS 2019 datasets. However, the integration of self-attention mechanisms increased computational complexity, leaving efficiency improvements necessary for large-scale data processing. Zhang, X. [14] designed the AM-DLFC model based on 2D CNN, integrating the CBAM module. This model demonstrated advantages in lightweight design and generalization across multiple datasets, yet remained susceptible to misclassifying specific actions and did not account for interference factors present in real-world scenarios.
@@ 5 Results & Discussion
In this section, we present a systematic and in-depth empirical analysis of the proposed USAD method. Specifically, we comprehensively evaluate the method from multiple perspectives, including the effectiveness of the model architecture, its adaptability to practical challenges, the impact of key mechanism parameters, and overall performance comparison.
@@ 5.2 Study on Addressing Class Imbalance
To systematically validate the effectiveness of the proposed USAD method in addressing class imbalance, experiments were conducted on three public datasets: WISDM (mild imbalance), PAMAP2 (moderate imbalance), and OPPORTUNITY (severe imbalance). The focus was on evaluating the impact of class imbalance on model performance, using both quantitative metrics and T-SNE visualization to compare the performance of USAD with various mainstream models. By benchmarking against classical CNN, recurrent neural networks (LSTM), basic residual networks (ResNet), attention-based models (SE-Res2Net), as well as high-performance existing models (MAG-Res2Net, ATFA[18], TCCSNet[19]), we assessed the strengths and weaknesses of each method from the perspectives of classification performance and feature space distribution.
@@ 5.2.1 Comparative Analysis of Classification Performance
Figure 4: Confusion matrices for WISDM and PAMAP2, highlighting USAD’s accurate classification of minority activities.
Figure 5: Confusion matrix for OPPORTUNITY, showing USAD’s strong generalization for complex rare behaviors.
@@
5.2.2 Analysis of Feature Discriminability
Model WISDM PAMAP2 OPPORTUNITY
Intra Inter Ratio Intra Inter Ratio Intra Inter Ratio
CNN 2.525 6.652 2.635 2.816 7.875 2.797 4.777 7.957 1.666
LSTM 1.047 6.242 5.961 1.726 6.883 3.988 2.043 6.449 3.156
ResNet50 0.446 6.508 14.594 1.583 6.930 4.379 3.225 7.774 2.410
SERes2Net 1.320 6.598 4.996 1.491 7.102 4.762 2.668 6.727 2.522
MAG-Res2Net 0.490 6.543 13.344 1.885 6.824 3.621 3.080 6.648 2.158
ATFA 2.797 6.930 2.478 3.007 9.907 3.294 4.850 8.379 1.728
TCCSNet 2.906 6.930 2.385 3.018 8.289 2.746 4.77 8.492 1.779
USAD 0.386 6.578 17.063 1.670 7.336 4.395 2.170 7.098 3.271
@@ 5.2.3 T-SNE Visualization Analysis
To more intuitively illustrate the feature space distributions of different models, we employed the T-SNE algorithm to reduce the dimensionality of feature vectors from the test samples and visualized them in a two-dimensional space. As shown in Figure 6 7 8, USAD exhibits outstanding feature separability across all three datasets. Except for a very small number of misclassified samples, USAD enables clear, tightly clustered, and well-separated distributions for different classes, effectively mitigating the adverse impact of class imbalance on the model’s discriminative ability.
Figure 6: T-SNE Visualization Comparison: Feature Distributions of CNN, ResNet50, TCCSNet, and USAD Models on the WISDM Dataset. Results of additional comparison models are provided in the Supplementary Material.
@@
Figure 7: T-SNE Visualization: Feature Distribution of CNN, ResNet50, TCCSNet, and USAD on the PAMAP2 Dataset. Results of additional comparison models are provided in the Supplementary Material.
@@
Figure 8: T-SNE Visualization: Feature Clustering of CNN, ResNet50, TCCSNet, and USAD on the OPPORTUNITY Dataset. Results of additional comparison models are provided in the Supplementary Material.
@@
5.3 Hyperparameter Exploration of the Adaptive Compound Loss Function on the PAMAP2 Dataset
5.3.1 Independent Optimization of Single Loss Function Hyperparameters
[…]
$\hat{y}$ Acc. (%) F1 (%) ECE
0 85.90 85.68 8.36
0.05 85.81 85.58 5.74
0.1 80.19 79.12 5.58
0.2 82.95 82.02 8.50
Table 4: Sensitivity analysis of label smoothing
coefficients.
$\gamma$ $\alpha$ Acc. (%) F1 (%)
1 0.25 92.69 92.50
2 0.25 91.68 91.35
3 0.25 89.74 88.62
1 0.5 91.84 91.67
1 0.75 92.38 92.15
Table 5: Focal loss hyperparameter tuning.
$\beta$ Acc. (%) F1 (%)
0.9 93.00 92.99
0.99 90.82 90.52
0.999 89.50 88.88
Table 6: Class-balanced loss parameter optimization.
5.3.2 Experiments on Composite Loss Weights and Regulation Parameters
[0.33, 0.33, 0.34] 1 0.1 0.8 93.16 93.06 92.03 91.85 Best Performance
[0.6, 0.2, 0.2] 1 0.1 0.8 84.19 83.78 83.96 85.60 Weight Bias
[0.2, 0.6, 0.2] 1 0.1 0.8 78.10 75.94 83.49 82.14 Weight Bias
[0.33, 0.33, 0.34] 0.5 0.1 0.8 79.81 76.83 81.12 80.26 Low Temperature
[0.33, 0.33, 0.34] 2 0.1 0.8 78.19 75.87 78.45 79.43 High Temperature
[0.33, 0.33, 0.34] 1 0.05 0.9 78.19 75.87 78.45 79.43 Wide Weight Range
Table 7: Performance Comparison with Different Hyperparameters
5.3.3 Sensitivity Analysis and Interleaved Experiments
$\hat{y}$ Accuracy (%) F1 Score (%)
0.0 86.19 85.68
0.05 75.33 73.17
0.1 86.67 86.66
0.2 78.29 77.13
Table 8: $\hat{y}$ Model Performance
Relationship
$\gamma$ $\alpha$ Accuracy (%) F1 Score (%)
1 0.25 90.86 90.75
2 0.25 82.48 80.68
3 0.25 77.43 74.38
1 0.5 91.05 90.65
1 0.75 87.71 87.37
Table 9: $\gamma$ and $\alpha$ Model Performance Relationship
$\beta$ Accuracy (%) F1 Score (%)
0.9 91.33 90.78
0.99 91.52 89.97
0.999 94.24 93.86
Table 10: $\beta$ Model Performance Relationship

### v2

@@ 2.1 Deep Feature Extraction
High-quality feature extraction is fundamental to accurate complex human activity classification and recognition. Recent research on neural network architecture design and multi-source HAR feature integration has advanced significantly, yet conventional ResNet architectures still suffer from weak inter-channel correlations, parameter redundancy, and insufficient feature reutilization.. To address these, Zhang et al.. [1] proposed ResNeSt, introducing a Split-Attention module and restructured feature aggregation to enhance channel dependency modeling and feature representation. Mekruksavanich et al. [2] adopted ResNeXt, integrating inertial and stretch sensor data for efficient activity classification. Other studies enrich feature extraction methods: Y. Wang et al. [3] proposed DMEFAM, integrating temporal-channel-spatial features via Bi-GRU, self-attention, CBAM, and ResNet-18, improving accuracy and DAAD generalization. Zhang J. et al. [4] used contrastive learning to extract high-precision activity-style features from multi-source data, preserving data characteristics and class semantics for enhanced cross-subject generalization. Liu X. et al. [5] encoded time series into RGB images via MTF/RP/GAF, designing CIR-DFENet with multi-scale convolution, global attention, and 1D CNN+LSTM for effective multi-source feature fusion.
@@ 2.2 Data Augmentation
A variety of data augmentation methods have been proposed to address data scarcity and model overfitting. Early work by J. Wang [6] introduced SensoryGANs, leveraging generative adversarial networks (GANs) to generate synthetic sensor data for improved activity recognition accuracy, alongside visualization methods for evaluating synthetic data quality. However, SensoryGANs faced challenges including unstable training dynamics and the requirement for activity-specific models, which significantly reduced computational efficiency. Li, X [7] proposed ActivityGAN to generate multi-activity data within a unified framework, effectively expanding datasets and enhancing model training efficacy and classification accuracy. Nevertheless, it failed to yield substantial improvements in recognition performance and exhibited notable limitations in cross-activity generalization capabilities.
To overcome these limitations of traditional approaches, researchers have turned to denoising diffusion probabilistic models (DDPM) [8]. S. Shao [9] designed a diffusion modeling framework tailored to sensor data characteristics, reconstructing the UNet architecture to handle low-channel and variable-length sequences, simplifying sampling layers to avoid over-feature extraction, and embedding class labels for conditional generation. Experimental results on HASC and PAMAP2 datasets demonstrated that the quality of generated data surpassed traditional transformation methods and GAN-based approaches, with a 1:4 real-synthetic data blending ratio further optimizing model performance. Current diffusion-based methods predominantly focus on unimodal data (e.g., inertial sensors), while research on collaborative augmentation for multimodal data (e.g., IMU sensors) in human activity recognition (HAR) tasks remains insufficient. Although Zhou, Y [10] adopted gradient optimization to determine augmentation strategy probability distributions, enabling automatic and efficient policy search—operating in two stages, accommodating multimodal properties, and demonstrating model-agnostic scalability—it exhibited limited performance gains in adjacent domains (e.g., EEG classification), highlighting that modality specificity inherently constrains the universality of augmentation strategies
@@ 2.3 Attention Mechanisms
In recent years, attention-based deep learning models have gained traction. Abedini, A. [11] proposed Att-CNN-BLSTM, using CNN and BLSTM to extract spatial-temporal features with an attention layer, achieving 90.54% accuracy and 86.88% F1-score, outperforming traditional models, but with high computational complexity and limited interpretability. Ye, X. et al. [12] developed DGDA, combining generative models with temporal attention to exploit time series dependencies, improving distribution alignment for HAR with excellent cross-user performance, yet limited in fusing heterogeneous sensors and resource-heavy. Zhao, J. et al. [13] integrated self-attention into CNN-BiLSTM for DDoS detection, achieving 95.67% accuracy via 1D CNN-BiLSTM feature extraction and attention emphasis, but with increased complexity needing efficiency optimization. Zhang, X. [14] designed AM-DLFC with 2D CNN and CBAM, lightweight and generalizable, but prone to misclassifying specific actions and ignoring real-world interferences.
@@ 4 Results & Discussion
In this section, we present a systematic and in-depth empirical analysis of the proposed USAD method. Specifically, we comprehensively evaluate the method from multiple perspectives, including the effectiveness of the model architecture, its adaptability to practical challenges, the impact of key mechanism parameters (Appendix C), and overall performance comparison.
@@ 4.2 Study on Addressing Class Imbalance
To systematically validate the effectiveness of the proposed USAD method in addressing class imbalance, experiments were conducted on three public datasets: WISDM (mild imbalance), PAMAP2 (moderate imbalance), and OPPORTUNITY (severe imbalance). The focus was on evaluating the impact of class imbalance on model performance, using both quantitative metrics (Appendix B) and T-SNE visualization to compare the performance of USAD with various mainstream models. By benchmarking against classical CNN, recurrent neural networks (LSTM), basic residual networks (ResNet), attention-based models (SE-Res2Net), as well as high-performance existing models (MAG-Res2Net, ATFA[15], TCCSNet[16]), we assessed the strengths and weaknesses of each method from the perspectives of classification performance and feature space distribution.
@@ 4.2.1 Comparative Analysis of Classification Performance
Figure 4: Confusion matrices for WISDM, PAMAP2, and OPPORTUNITY.
@@ 4.2.2 T-SNE Visualization Analysis
To more intuitively illustrate the feature space distributions of different models, we employed the T-SNE algorithm to reduce the dimensionality of feature vectors from the test samples and visualized them in a two-dimensional space. As shown in Figure 5, USAD exhibits outstanding feature separability across all three datasets. Except for a very small number of misclassified samples, USAD enables clear, tightly clustered, and well-separated distributions for different classes, effectively mitigating the adverse impact of class imbalance on the model’s discriminative ability.
Figure 5: T-SNE Visualization Comparison: Feature Distributions of CNN, ResNet50, TCCSNet, and USAD Models on the (a) WISDM, (b) PAMAP2, and (c) OPPORTUNITY Dataset. Results of additional comparison models are provided in the Supplementary Material.
@@ 5 Conclusion
Appendix A Experimental Design
A.1 Datasets
A.2 Evaluation Metrics
[…]
Appendix B Analysis of Feature Discriminability
Model WISDM PAMAP2 OPPORTUNITY
Intra Inter Ratio Intra Inter Ratio Intra Inter Ratio
CNN 2.525 6.652 2.635 2.816 7.875 2.797 4.777 7.957 1.666
LSTM 1.047 6.242 5.961 1.726 6.883 3.988 2.043 6.449 3.156
ResNet50 0.446 6.508 14.594 1.583 6.930 4.379 3.225 7.774 2.410
SERes2Net 1.320 6.598 4.996 1.491 7.102 4.762 2.668 6.727 2.522
MAG-Res2Net 0.490 6.543 13.344 1.885 6.824 3.621 3.080 6.648 2.158
ATFA 2.797 6.930 2.478 3.007 9.907 3.294 4.850 8.379 1.728
TCCSNet 2.906 6.930 2.385 3.018 8.289 2.746 4.77 8.492 1.779
USAD 0.386 6.578 17.063 1.670 7.336 4.395 2.170 7.098 3.271
Appendix C Hyperparameter Exploration of the Adaptive Compound Loss Function on the PAMAP2 Dataset
C.1 Independent Optimization of Single Loss Function Hyperparameters
[…]
$\hat{y}$ Acc. (%) F1 (%) ECE
0 85.90 85.68 8.36
0.05 85.81 85.58 5.74
0.1 80.19 79.12 5.58
0.2 82.95 82.02 8.50
Table 7: Sensitivity analysis of label smoothing
coefficients.
$\gamma$ $\alpha$ Acc. (%) F1 (%)
1 0.25 92.69 92.50
2 0.25 91.68 91.35
3 0.25 89.74 88.62
1 0.5 91.84 91.67
1 0.75 92.38 92.15
Table 8: Focal loss hyperparameter tuning.
$\beta$ Acc. (%) F1 (%)
0.9 93.00 92.99
0.99 90.82 90.52
0.999 89.50 88.88
Table 9: Class-balanced loss parameter optimization.
C.2 Experiments on Composite Loss Weights and Regulation Parameters
[0.33, 0.33, 0.34] 1 0.1 0.8 93.16 93.06 92.03 91.85 Best Performance
[0.6, 0.2, 0.2] 1 0.1 0.8 84.19 83.78 83.96 85.60 Weight Bias
[0.2, 0.6, 0.2] 1 0.1 0.8 78.10 75.94 83.49 82.14 Weight Bias
[0.33, 0.33, 0.34] 0.5 0.1 0.8 79.81 76.83 81.12 80.26 Low Temperature
[0.33, 0.33, 0.34] 2 0.1 0.8 78.19 75.87 78.45 79.43 High Temperature
[0.33, 0.33, 0.34] 1 0.05 0.9 78.19 75.87 78.45 79.43 Wide Weight Range
Table 10: Performance Comparison with Different Hyperparameters
C.3 Sensitivity Analysis and Interleaved Experiments
$\hat{y}$ Accuracy (%) F1 Score (%)
0.0 86.19 85.68
0.05 75.33 73.17
0.1 86.67 86.66
0.2 78.29 77.13
Table 11: $\hat{y}$ Model Performance
Relationship
$\gamma$ $\alpha$ Accuracy (%) F1 Score (%)
1 0.25 90.86 90.75
2 0.25 82.48 80.68
3 0.25 77.43 74.38
1 0.5 91.05 90.65
1 0.75 87.71 87.37
Table 12: $\gamma$ and $\alpha$ Model Performance Relationship
$\beta$ Accuracy (%) F1 Score (%)
0.9 91.33 90.78
0.99 91.52 89.97
0.999 94.24 93.86
Table 13: $\beta$ Model Performance Relationship